The swordfish is the fastest of all fish. If it accelerates at a rate of 4.0 m / s2 for 1.7 s from its initial speed of 42 km / h, what is its final speed?


1
Expert's answer
2021-01-06T13:53:44-0500

Solution

According to equation of motion is given by

"v=u+at"

Initial velocity u=42km/h=11.67m/s

Final velocity

"v=11.67+4\times1.7=18.47m/s"